OpenStudy (anonymous):

Solve C = 3(2f – 5h) for f in terms of C and h. You must show all work.

mathslover (mathslover):

\[\huge{6f-15h=C}\] \[\huge{6f=C+15h}\] \[\huge{f=\frac{C+15h}{6}}\]

mathslover (mathslover):

u must know that \(\huge{a(b+c)=ab+ac}\)

OpenStudy (anonymous):

Thank you very much.